OpenStudy (anonymous):

An object with a kinetic energy of 2160J has a mass of 120kg. What is its velocity?

OpenStudy (anonymous):

What is the equation for kinetic energy?

OpenStudy (anonymous):

ke=1/2mv2

OpenStudy (anonymous):

So you have the KE, the mass. The only unknown left is v.

OpenStudy (anonymous):

could you solve it

OpenStudy (anonymous):

please

OpenStudy (stamp):

\[KE=1/2 \ mv^2\] KE = 2160 J m = 120 kg KE = 1/2 mv^2 2KE = mv^2 2KE / m = v^2 v = sqrt( 2KE / m ) v = sqrt( 2(2160) / 120 ) v = sqrt( 4320 / 120 ) v = 36

OpenStudy (stamp):

Sometimes in Physics calculations, it is easier to solve for your variable and then substitute values, rather than put in values and then juggle all the numbers.